# Online Flow Time Minimization: Tight Bounds for Non-Preemptive Algorithms

### v2

@@ Abstract
This paper studies the classical online scheduling problem of minimizing total flow time for $n$ jobs on $m$ identical machines. Prior work often cites the $\Omega(n)$ lower bound for non-preemptive algorithms to argue for the necessity of preemption or resource augmentation, which shows the trivial $O(n)$-competitive greedy algorithm is tight. However, this lower bound applies only to deterministic algorithms in the single-machine case, leaving several fundamental questions unanswered. Can randomness help in the non-preemptive setting, and what is the optimal online deterministic algorithm when $m\geq 2$? We resolve both questions. We present a polynomial-time randomized algorithm with competitive ratio $\Theta(\sqrt{n/m})$ and prove a matching randomized lower bound, settling the randomized non-preemptive setting for every $m$. This also improves the best-known offline approximation ratio from $O(\sqrt{n/m}\log(n/m))$ to $O(\sqrt{n/m})$. On the deterministic side, we present a non-preemptive algorithm with competitive ratio $O(n/m^{2}+\sqrt{n/m}\log m)$ and prove a nearly matching lower bound.
Our framework also extends to the kill-and-restart model, where we reveal a sharp transition of deterministic algorithms: we design an asymptotically optimal algorithm with the competitive ratio $O(\sqrt{n/m})$ for $m\geq 2$, yet establish a strong $\Omega(n/\log n)$ lower bound for $m=1$. Moreover, we show that randomization provides no further advantage, as the lower bound coincides with that of the non-preemptive setting.
While our main results assume prior knowledge of $n$, we also investigate the setting where $n$ is unknown. We show kill-and-restart is powerful enough to break the $O(n)$ barrier for $m\geq 2$ even without knowing $n$. Conversely, we prove randomization alone is insufficient, as no algorithm can achieve an $o(n)$ competitive ratio in this setting.
@@ 1 Introduction
The scheduling problem is a core combinatorial optimization problem across various disciplines in computer science. Its significance stems not only from its wide range of applications but also from its simple yet representative formulation. Numerous variants of scheduling problems exist, motivated by diverse real-world scenarios, such as different machine types (identical, related, or unrelated) and varying objectives (makespan, completion time, flow time). A prominent line of research in scheduling focuses on making irrevocable decisions for jobs that arrive online over time — a framework known as online scheduling, which was initiated in the 1960s by graham1966bounds. Consider an online service platform that must dispatch jobs submitted by multiple users as they arrive; this naturally leads to the objective of minimizing the total flow time, the sum over jobs of their completion time minus release time.
In this paper, we focus on the flow time minimization problem on $m$ identical machines. Most studies on the flow time objective allow additional capabilities for algorithms, such as preemption, resource augmentation, and rejection. Many papers (e.g., [DBLP:conf/focs/GargK07, DBLP:conf/esa/LucarelliTST16, DBLP:conf/stoc/ChadhaGKM09, DBLP:conf/soda/ChenIP25, DBLP:conf/icalp/GuptaKL18, DBLP:journals/dam/EpsteinS06]) have argued that they are necessary to surpass the trivial greedy algorithm [^1] The trivial greedy algorithm simply schedules any waiting job whenever any machine becomes available, achieving a competitive ratio of $\Theta(n/m)$. because of the following single-machine lower bound consisting of $n$ jobs: the adversary releases a unit job at time 0 and, upon scheduling, reveals n1n-1 $\varepsilon$-jobs, forcing any deterministic non-preemptive algorithm to incur $\Omega(n)$ flow time while the offline optimum remains $O(1)$ [DBLP:journals/siamcomp/KellererTW99, DBLP:conf/stoc/ChekuriKZ01]. However, this lower bound applies only to deterministic algorithms in the single-machine setting. The limitations of this lower bound motivate our exploration of the following open questions.
@@
Randomization is widely employed in the design of online algorithms. Against an oblivious adversary who does not adapt to the randomness used by the algorithm, randomization often breaks deterministic barriers, including the classic caching problem and matching problem [DBLP:journals/jal/FiatKLMSY91, DBLP:conf/stoc/KarpVV90].
In our problem, the previously known lower bound $\Omega(\sqrt{n})$ for oblivious adversaries by DBLP:journals/tcs/EpsteinS03 applies only to single-machine; no information-theoretic lower bound stronger than $\Omega(1)$ was previously known when $m\geq 2$. A gap remains between the current best-known $\Theta(n/m)$-competitive trivial greedy algorithm and the largely unanswered questions about lower bounds.
Question for Randomized Algorithms Can we use randomization to beat the trivial greedy algorithm? Is it possible to find an algorithm much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
@@
In fact, even without randomization, the problem remains unsolved when $m$ is large: the deterministic lower bound $\Omega(n/m^{2})$ established by DBLP:journals/dam/EpsteinS06 decreases much faster than the trivial upper bound $\Theta(n/m)$ achieved by the greedy algorithm, leaving a substantial gap when $m$ is comparable to $n$.
Question for Deterministic Algorithms Is greedy indeed the best deterministic algorithm for multi-machine case? What is the optimal competitive ratio when $m$ is large?
@@
Another direction that is closely related to the non-preemptive setting is the kill-and-restart capability, first introduced by DBLP:journals/siamcomp/ShmoysWW95, where the scheduler is allowed to kill a job and restart it from scratch, losing all processed work. Kill-and-restart lies between preemption and non-preemption. This model is particularly suitable in settings where partially processed work cannot be resumed due to system or application constraints (e.g., data corruption or non-checkpointable tasks). It also has a practical advantage over preemption, as it avoids the need to maintain and store potentially unbounded intermediate states of suspended jobs. This notion has been studied under objectives such as makespan and total completion time, where it has been shown to surpass previously known lower bounds [DBLP:conf/approx/0002KTWZ18, DBLP:journals/mp/JagerSWW25, DBLP:journals/jal/SteeP05]. Notably, DBLP:journals/mp/JagerSWW25 considered online non-clairvoyant scheduling with the objective of minimizing total completion time and demonstrated that kill-and-restart significantly improves the competitive ratio from $\Omega(n)$ to $O(1)$.
Unlike the completion time objective where kill-and-restart has proven powerful, its potential for flow time minimization has remained largely unexplored. No algorithms were previously known to utilize the kill-and-restart capability in the flow time minimization problem, and the only known lower bound $\Omega(\sqrt{n})$ by DBLP:journals/tcs/EpsteinS03 for the kill-and-restart model is also restricted to the single-machine case. Since the only known algorithm without preemption is the $O(n/m)$-competitive greedy algorithm, it remained unclear whether the kill-and-restart capability can improve competitiveness in flow-time minimization, and to what extent such improvement is possible.
Question for Kill-and-Restart Can we use kill-and-restart to beat the trivial greedy algorithm? Is it possible to find an algorithm much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
1.1 Our Results
We give a complete set of answers to the above questions, summarized in Table 1 (the single-machine case) and Table 2 (the multi-machine case). A key technical feature of our lower bounds in the multi-machine case is their robustness: unlike bounds that only apply to specific choices of these parameters, they hold for every fixed pair of $(n,m)$.
The results presented in this section assume prior knowledge of $n$. This assumption is common when we study the non-preemptive version of the flow time minimization problem [DBLP:journals/dam/EpsteinS06, DBLP:journals/tcs/EpsteinS03], and we will further discuss this in Section 1.2 and present results on the algorithms without prior knowledge of $n$.
@@ 1.1 Our Results
We first address whether randomization can overcome the $\Omega(n)$ deterministic barrier for non-preemptive algorithms. We answer this affirmatively by presenting an $O(\sqrt{n})$-competitive randomized algorithm for the single-machine case, matching the established lower bound by DBLP:journals/tcs/EpsteinS03. We then extend this to the multi-machine setting, achieving an $O(\sqrt{n/m})$ competitive ratio. This result is optimal, as we prove a matching $\Omega(\sqrt{n/m})$ lower bound that holds for every pair of $(n,m)$. This completely characterizes the power of randomization in the non-preemptive setting.
@@
Perhaps our most surprising finding is non-uniform impact of the kill-and-restart capability on deterministic algorithms.
  • For the single-machine case, we show that kill-and-restart offers little advantage, proving a strong lower bound of $\Omega(n/\log n)$.
  • However, with just one additional machine, the problem’s competitive ratio collapses. We present a deterministic algorithm with kill-and-restart that achieves an $O(\sqrt{n/m})$ competitive ratio for all $m\geq 2$. This is a quadratic improvement over the single-machine case and remarkably matches the optimal bound achievable with randomization.
Conversely, we show that for randomized algorithms, kill-and-restart provides no additional power, as the $\Omega(\sqrt{n/m})$ lower bound holds even in this stronger model.
@@
A direct consequence of our online results is a new state-of-the-art for the classic offline problem. We improve the long-standing best approximation ratio from $O(\sqrt{n/m}\log(n/m))$ by DBLP:journals/jcss/LeonardiR07 to $O(\sqrt{n/m})$. The improved bound can be realized by either derandomizing our randomized algorithm or applying our deterministic kill-and-restart algorithm in an offline context where restarts can be ignored at no cost.
@@
[DBLP:journals/siamcomp/KellererTW99, DBLP:conf/stoc/ChekuriKZ01] Theorem 7
@@
[DBLP:journals/tcs/EpsteinS03] [DBLP:journals/tcs/EpsteinS03]
@@
1.2 Further Discussion of the Prior Knowledge of $n$
In our main results, we follow the common convention in non-preemptive flow-time scheduling by assuming that $n$, the total number of jobs, is known in advance. This assumption is common to nearly all significant prior work in this area, including both upper and lower bounds [DBLP:journals/dam/EpsteinS06, DBLP:journals/tcs/EpsteinS03]. The necessity of this assumption is underscored by strong lower bounds showing that without it, no deterministic algorithm can beat the trivial $O(n)$ competitive ratio, even if the algorithm can use $\ell$ machines but the optimal solution can only use one [DBLP:journals/dam/EpsteinS06].
However, this barrier does not preclude the possibility that randomization or the kill-and-restart capability could yield improvements in the more challenging “unknown nn” setting. We investigate this question and show that these two tools have different capabilities:
  • Randomization alone is insufficient. We prove that for a single machine, no randomized algorithm can achieve a competitive ratio of $o(n)$ without knowing $n$. For the multi-machine setting, we further show that no randomized algorithm can achieve a competitive ratio of $o(n/m^{2}+\sqrt{n/m})$ without prior knowledge of $n$. This demonstrates that knowing $n$ is necessary for designing non-trivial randomized algorithms in the non-preemptive setting.
  • Kill-and-restart overcomes the barrier. In sharp contrast, we show that the kill-and-restart model is powerful enough to break the $O(n)$ barrier when $m\geq 2$. We extend our deterministic kill-and-restart algorithm to operate without knowledge of $n$ and prove that it achieves an $O(n^{\alpha}/\sqrt{m})$-competitive ratio, where $\alpha=(\sqrt{5}-1)/2\approx 0.618$ is the golden ratio. This demonstrates that the kill-and-restart capability provides a fundamental advantage over purely randomized approaches in this setting.
Finally, we note that our algorithms are robust to estimation. If only an approximate estimate $\hat{n}\in[n/c,cn]$ is known instead of the accurate value of $n$, as might be provided by heuristics or machine-learned advice, all algorithms’ guarantees hold within a constant factor of $c$, preserving their asymptotic performance.
@@ 1.3 Our Techniques
A central challenge in non-preemptive scheduling is to partition incoming jobs into two manageable groups: a small set of large jobs, and a set of small jobs whose processing times are provably limited. This distinction is key to circumventing the classic $\Omega(n)$ lower bound, where handling a single large job at the wrong time can delay a swarm of subsequent small jobs, leading to high flow time. Offline algorithms can address this by partitioning jobs based on global properties of the total job set (e.g., the offline approximation algorithms by DBLP:journals/siamcomp/KellererTW99 and DBLP:journals/jcss/LeonardiR07 relied on a pre-solved preemptive solution over all jobs, and use the flow time of jobs in the preemptive solution to partition jobs), but this is impossible to implement in an online context, since future jobs are unknown. Therefore, a fundamentally new approach is required for the online setting where decisions are irrevocable and based on incomplete information.
@@
  • Managing dynamic re-classification. A job initially classified as large may later become small as larger jobs arrive. If we start treating it as a small job from that point on, we must carefully bound its previous waiting time and analyze whether this transition delays future jobs by effectively leaving one more job behind OPT. To handle this transition gracefully, we introduce the concept of proxy jobs. Instead of re-label a large job when it should become small, we treat it as a newly arrived small job, called a proxy job. We then decompose its total flow time into two parts: the waiting time during its large phase, and the flow time of the proxy job. Note that a proxy job is not an actual job from the original instance, so we must further bound the additional impact introduced by creating such proxy jobs.
@@
With this online rank-based partitioning framework in place, subsequent algorithmic design can focus solely on coordinating a bounded number of large jobs with an efficiently scheduled set of small jobs. A key advantage of this framework is that it no longer requires a preemptive algorithm to exploit global properties of the entire job set. This distinction not only enables our algorithm to be implemented online, but also improves the offline approximation ratio by a logarithmic factor over the result of DBLP:journals/jcss/LeonardiR07, since their logarithmic loss arises from relying on a suboptimal preemptive algorithm in the multi-machine setting.
We believe that our online partitioning framework and the analysis of NSJF provide a general technical insight that can be applied to broader classes of non-preemptive scheduling problems, such as the flow time minimization problem in the related machine setting.

### v3

@@ Abstract
This paper studies the online scheduling problem of minimizing total flow time for $n$ jobs on $m$ identical machines. A classical $\Omega(n)$ lower bound shows that no deterministic single-machine algorithm can beat the trivial greedy, even when $n$ is known in advance. However, this barrier is specific to deterministic algorithms on a single machine, leaving open what randomization, multiple machines, or the kill-and-restart capability can achieve.
We give a nearly complete answer. For randomized non-preemptive algorithms, we establish a tight $\Theta(\sqrt{n/m})$ competitive ratio, which also improves the best offline approximation to $O(\sqrt{n/m})$. For deterministic non-preemptive algorithms on multiple machines, we prove an $O(n/m^{2}+\sqrt{n/m}\log m)$ upper bound and an $\Omega(n/m^{2}+\sqrt{n/m})$ lower bound. In the kill-and-restart model, we reveal a sharp transition for deterministic algorithms: $\Omega(n/\log n)$ for $m=1$ versus $\Theta(\sqrt{n/m})$ for $m\geq 2$; the latter matches the optimal randomized ratio, and we further show that randomization provides no additional power in this model.
We also investigate the setting where $n$ is unknown. We prove that no randomized non-preemptive algorithm achieves $o(n)$ on one machine or $o(n/m^{2}+\sqrt{n/m})$ on $m$ machines. In contrast, our kill-and-restart algorithm achieves $O(n^{\alpha}/\sqrt{m})$ for $m\geq 2$, where $\alpha=(\sqrt{5}-1)/2$, breaking the trivial bound without knowledge of $n$.
@@ 1 Introduction
Scheduling is one of the most fundamental problems in combinatorial optimization, with deep roots in both theory and practice. A rich landscape of scheduling variants spans different machine environments (identical, related, and unrelated [8, 24, 13]) and optimization objectives (makespan [15], completion time [18], and flow time [2, 23, 25]). A prominent line of research focuses on the online setting [16, 3], where jobs arrive over time and the scheduler must make irrevocable decisions without knowledge of future arrivals. In modern applications such as cloud resource management and data center scheduling [21, 14, 12], an online service platform must dispatch jobs submitted by multiple users as they arrive; this naturally leads to the objective of minimizing the total flow time (or equivalently, the average flow time), defined as the sum over all jobs of the difference between completion time and release time.
In this paper, we focus on the flow time minimization problem on $m$ identical machines. Most positive results (e.g., [27, 4, 6, 17, 10]) on the flow time objective require additional capabilities — such as preemption, resource augmentation, and rejection — to surpass the trivial greedy algorithm [^1] The trivial greedy algorithm simply schedules any waiting job whenever any machine becomes available, achieving a competitive ratio of $\Theta(n/m)$.. This is due to the following single-machine lower bound, which holds even when $n$ is known in advance:
The adversary first releases a unit job at time 0. As soon as the algorithm starts processing it or after waiting for $n$ units of time, the adversary releases n1n-1 $\varepsilon$-jobs.
Any deterministic non-preemptive algorithm is thus forced to incur $\Omega(n)$ flow time, whereas the offline optimum remains $O(1)$ [23, 5]. However, this lower bound applies only to deterministic algorithms in the single-machine setting. The limitations of this lower bound motivate our exploration of the following open questions.
@@
Randomization is widely employed in the design of online algorithms. Against an oblivious adversary who does not adapt to the randomness used by the algorithm, randomization often breaks deterministic barriers, including the classic caching problem and matching problem [11, 22].
In our problem, the previously known lower bound $\Omega(\sqrt{n})$ for oblivious adversaries by [9] applies only to the single-machine case; for $m\geq 2$, no information-theoretic lower bound stronger than $\Omega(1)$ was previously known, leaving a substantial gap from the $O(n/m)$ competitive ratio of the trivial greedy algorithm.
Open Question 1: Randomized Algorithms Can we use randomization to beat the trivial greedy algorithm? Is it possible to find a randomized algorithm much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
@@
In fact, even without randomization, the problem remains unsolved when $m$ is large: the deterministic lower bound $\Omega(n/m^{2})$ established by [10] decreases much faster than the trivial upper bound $\Theta(n/m)$ achieved by the greedy algorithm, leaving a substantial gap when $m$ is comparable to $n$.
Open Question 2: Deterministic Algorithms Is greedy indeed the best deterministic algorithm for the multi-machine case? What is the optimal competitive ratio when $m$ is large?
@@
A closely related direction is the kill-and-restart model, first introduced by [29], where the scheduler may kill a running job and restart it from scratch, losing all processed work. This model is particularly well-suited to applications where partially processed work cannot be resumed (e.g., data corruption or non-checkpointable tasks). Unlike preemption, it avoids the need to maintain and store potentially unbounded intermediate states of suspended jobs, preserving the non-preemptive nature of job execution while giving the scheduler additional flexibility to correct past decisions.
Kill-and-restart has proven powerful for other scheduling objectives. For makespan and total completion time, it surpasses known non-preemptive barriers [19, 30]; most notably, [20] showed that for non-clairvoyant scheduling under total completion time, kill-and-restart improves the competitive ratio from $\Omega(n)$ to $O(1)$.
Despite these successes, the potential of kill-and-restart for flow time minimization has remained largely unexplored. No algorithm was previously known to exploit this capability, and the only known lower bound $\Omega(\sqrt{n})$ [9] is restricted to the single-machine case. It thus remained unclear whether, and to what extent, kill-and-restart can improve upon the $O(n/m)$-competitive greedy algorithm.
Open Question 3: Kill-and-Restart Can we use kill-and-restart capability to beat the trivial greedy algorithm? Is it possible to find an algorithm under the kill-and-restart model that is much better than $\Theta(\sqrt{n})$ (perhaps even a constant) when $m\geq 2$?
1.1 Our Results with Known $n$
We give a complete set of answers to Open Questions 1–3 when $n$ is known in advance, summarized in Table 1 (the single-machine case) and Table 2 (the multi-machine case). A key technical feature of our lower bounds in the multi-machine case is their robustness: unlike bounds that only apply to specific choices of these parameters, they hold for every fixed pair of $(n,m)$.
Throughout this paper, we assume $n\geq m$ when presenting competitive ratio; when $n<m$, all jobs can be processed immediately without waiting, yielding optimal schedules with a competitive ratio of $1$.
@@ 1.1 Our Results with Known $n$
We show that randomization can overcome the $\Omega(n)$ deterministic barrier. For the single-machine case, we present an $O(\sqrt{n})$-competitive randomized algorithm, matching the lower bound of [9]. For the multi-machine case, we achieve an $O(\sqrt{n/m})$ competitive ratio and prove a matching $\Omega(\sqrt{n/m})$ lower bound that holds for every pair $(n,m)$. Together, these results completely characterize the power of randomization in non-preemptive flow time minimization.
@@
Perhaps our most surprising finding is the non-uniform impact of the kill-and-restart capability on deterministic algorithms.
  • For the single-machine case, kill-and-restart offers little advantage: we prove a lower bound of $\Omega(n/\log n)$.
  • For $m\geq 2$, the competitive ratio drops dramatically. We present a deterministic kill-and-restart algorithm achieving $O(\sqrt{n/m})$, a quadratic improvement over the single-machine case.
Remarkably, this deterministic kill-and-restart ratio of $O(\sqrt{n/m})$ matches the optimal randomized bound. We further show that this coincidence is tight: the $\Omega(\sqrt{n/m})$ lower bound holds even for randomized algorithms with kill-and-restart.
@@
A direct consequence of our online results is a new state-of-the-art for the classic offline problem. We improve the long-standing best approximation ratio from $O(\sqrt{n/m}\log(n/m))$ by [25] to $O(\sqrt{n/m})$. The improved bound can be realized by either derandomizing our randomized algorithm or applying our deterministic kill-and-restart algorithm in an offline context where restarts can be ignored at no cost.
@@
Shaded cells indicate results from this paper.
@@
[23, 5] Theorem 7
@@
[9] [9]
@@
Shaded cells indicate results from this paper.
@@
1.2 Our Results without Prior Knowledge of $n$
All results above assume that the total number of jobs $n$ is known in advance, a convention shared by prior positive results on non-preemptive flow time minimization [10, 9]. This assumption is critical for deterministic algorithms: [10] showed that without knowledge of $n$, no deterministic algorithm on any constant number of machines can achieve a competitive ratio of $o(n)$, even when compared against a single-machine offline optimum. Is this assumption also critical for randomized algorithms and algorithms with kill-and-restart capability?
Open Question 4: Unknown Number of Jobs Can randomization or kill-and-restart still overcome the deterministic $\Theta(n)$ barrier when $n$ is not known in advance?
We find a sharp separation between the two.
Randomization is insufficient.
We prove that without knowledge of $n$, no randomized non-preemptive algorithm can achieve a competitive ratio of $o(n)$ on a single machine, or $o(n/m^{2}+\sqrt{n/m})$ on $m$ machines.
Kill-and-restart overcomes the barrier.
In contrast, we extend our deterministic kill-and-restart algorithm to operate without knowledge of $n$, achieving a competitive ratio of $O(n^{\alpha}/\sqrt{m})$ for all $m\geq 2$, where $\alpha=(\sqrt{5}-1)/2\approx 0.618$. This significantly improves over the trivial $O(n/m)$ bound and demonstrates that kill-and-restart provides a fundamental advantage that persists even without knowledge of $n$.
Algorithms with predicted $n$.
Finally, we note that if an approximate prediction $\hat{n}\in[n/c,cn]$ is available instead of the exact value of $n$, all our known-$n$ algorithms’ guarantees degrade by at most a constant factor of $c$, preserving their asymptotic performance.
@@ 1.3 Our Techniques
A central challenge in non-preemptive scheduling is to partition incoming jobs into two manageable groups: a small set of large jobs, and a set of small jobs whose processing times are provably limited. This distinction is key to circumventing the classic $\Omega(n)$ lower bound, where handling a single large job at the wrong time can delay a swarm of subsequent small jobs, leading to high flow time. Offline algorithms can address this by partitioning jobs based on global properties of the total job set (e.g., the offline approximation algorithms by [23] and [25] relied on a pre-solved preemptive solution over all jobs, and used the flow time of jobs in the preemptive solution to partition jobs), but this is impossible to implement in an online context, since future jobs are unknown. Therefore, a fundamentally new approach is required for the online setting where decisions are irrevocable and based on incomplete information.
@@
  • (1)
    Managing dynamic re-classification. A job initially classified as large may later become small as larger jobs arrive. If we start treating it as a small job from that point on, we must carefully bound its previous waiting time and analyze whether this transition delays future jobs by effectively leaving one more job behind OPT. To handle this transition gracefully, we introduce the concept of proxy jobs. Instead of relabeling a large job when it should become small, we treat it as a newly arrived small job, called a proxy job. We then decompose its total flow time into two parts: the waiting time during its large phase, and the flow time of the proxy job. Note that a proxy job is not an actual job from the original instance, so we must further bound the additional impact introduced by creating such proxy jobs.
  • (2)
@@
With this online rank-based partitioning framework in place, subsequent algorithmic design can focus solely on coordinating a bounded number of large jobs with an efficiently scheduled set of small jobs. A key advantage of this framework is that it no longer requires a preemptive algorithm to exploit global properties of the entire job set. This distinction not only enables our algorithm to be implemented online, but also improves the offline approximation ratio by a logarithmic factor over the result of [25], since their logarithmic loss arises from relying on a suboptimal preemptive algorithm in the multi-machine setting.
We remark that our online partitioning framework and the analysis of NSJF do not inherently rely on the machines being identical, and we expect them to be useful for broader settings such as related machines.